OpenStudy (anonymous):

Parallelogram ABCD has coordinates of A (1,5) B (6,3) C(3,-1) and D(-2,1). What are the coordinates of E, the intersection of diagonals AC and BD

OpenStudy (anonymous):

You can solve it by making equations of AC and BD and finding the point of intersection of the two equations by linear equations in two variables.

OpenStudy (anonymous):

@christtinaaaxo do you know how to find the equation of a straight line???

OpenStudy (anonymous):

please help me i don't know how to do it

OpenStudy (anonymous):

okay then, this should be easy:- Suppose I have two points (a,b) and (c,d) and I want an equation which passes through the two points so this is what I do:- (y-b)/(x-a) = (b-d)/ (a-c)

OpenStudy (anonymous):

Now, you make the equations of AC and BD.

OpenStudy (anonymous):

do you have to do y=mx+b

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is the formula for equation passing through two points..

OpenStudy (anonymous):

okay i'm sorry as you both can there is no hope for me and i will be failing my regent

OpenStudy (anonymous):

No you won't! Even if you would, solving it is worth a try.

OpenStudy (anonymous):

Go on, you have the points, I have time (as you see), put the values and give me two equations. It's very easy. Just try.

OpenStudy (anonymous):

do you have to do the midpoint formula first?

OpenStudy (anonymous):

No. I gave you an equation. You can use that for AC and BD. THAT is step 1. Step 2 is solving the values for x and y. See, the value which you will get will be common for the likes AC and BD to intersect. THAT is your answer.

OpenStudy (anonymous):

i got (2,2) as an answer and that is one of my choices

OpenStudy (anonymous):

there you go.

OpenStudy (anonymous):

thank you very much !

OpenStudy (anonymous):

No problemo. ;)